OpenStudy (anonymous):

determine the intervals of increasing and decreasing, as well as all critical points for h(x)=x^4-4x^3+10. give the critical points in x-y coords.

OpenStudy (mathteacher1729):

http://tutorial.math.lamar.edu/Classes/CalcI/ShapeofGraphPtI.aspx This is a super duper tutorial on how to graph a function without using a calculator using 1st and 2nd derivatives. :)

OpenStudy (anonymous):

h(x)=x^4-4x^3+10 h'(x)=4x^3-12x^2 set it = to 0 and solve

OpenStudy (anonymous):

i got 3

OpenStudy (anonymous):

it was after that step i don't know what to do

OpenStudy (anonymous):

x=0,3

OpenStudy (anonymous):

check x<0 check 0<x<3 check x>3

OpenStudy (anonymous):

for x=-1 4x^3-12x^2 -4-12=-16 negative number so function is decreasing x<0

OpenStudy (anonymous):

-,-,+????

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

you told me to check the values at the certain places and i got a -/-/+

OpenStudy (anonymous):

dec/dec/inc.... would be another way of saying it

OpenStudy (anonymous):

thank you for that... and the critical points would be 3 and 0? and i'm not sure how to give them in xy coords

OpenStudy (anonymous):

(3,0) (0,0)

OpenStudy (anonymous):

on that graph though... those aren't the spots they change concavity or whatever else... its' more like (3,-3) (0,10)

OpenStudy (phi):

by x,y coords, they mean x,h(x) coords i.e. y is just h(x) when you plot this function

OpenStudy (anonymous):

so are the points 3,0 and 0,0

OpenStudy (phi):

you want h(3) and h(0) for the second coord

OpenStudy (anonymous):

thanks